X/a+x/b=c; where a ≠ 0,b ≠0, a ≠ -b,solve for x
Answer by stanbon(75887)

X/a+x/b=c; where a ≠ 0,b ≠0, a ≠ -b,solve for x
-----------------------
bx + ax = abc
x(b+a) = abc
------
x = abc/(a+b)
======================
Cheers,
Stan H.
============